OpenStudy (anonymous):

integrate from 0 to 2 ln (x+1)...... I used integration by parts but I am stuck after setting up the new integral

terenzreignz (terenzreignz):

\[\large \int_0^2 \ln(x+1)dx\] This it?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well, actually, there was no need for integration by parts... Can you evaluate this antiderivative? \[\large \int \ln \ x \ dx\]

OpenStudy (anonymous):

There's no need for integrating by parts... let u = x + 1. Now we have \(\int_1^3\ln u \ \mathrm{d}u\).

OpenStudy (anonymous):

the homework assignment asks to use parts

OpenStudy (anonymous):

... though, if you don't know \(\int\ln x\ \mathrm{d}x\) then you can determine it using integration by parts. Let \(u=\ln x\), \(\mathrm{d}v=\mathrm{d}x\) and thus \(\mathrm{d}u=\frac1x\mathrm{d}x\), \(v=x\). Now we have the integral \(\int u\ \mathrm{d}v=uv-\int v\ \mathrm{d}u\). Can you reduce now?

OpenStudy (anonymous):

ok ill give it a shot thanks for your help

terenzreignz (terenzreignz):

Must have been confused.... yeah, integration by parts was key here, my bad...

OpenStudy (anonymous):

Not really, @terenzreignz. Typically you'd already know that integral.

terenzreignz (terenzreignz):

Still, no way to derive it other than through integration by parts. There's no escaping it. And no, we were not taught that integral before we learned integration by parts, just like we were not taught the integral of tan x before we derived the integral of 1/x, etc...

OpenStudy (anonymous):

Yes I agree with you terenzreignz.

OpenStudy (anonymous):

@ oldrin.bataku It seemed too easy to just do a simple u sub, thus being in chap 7 integration by parts i kept forcing myself to use that method to solve the problem. Yet I am still stuck after setting up the equation for parts.